MIDC/Data Quality Statements
Total Records: 18, displaying 1-18

R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSDiffuse PSP03/06/201415:00:00Eppley AC ventilator fan has failed. Until fan can be replaced the fuse was removed so that it does not generate heat. PSP is running un-ventilated.
2BMSDiffuse PSP02/18/201413:00:0003/06/201415:00:00Instrument was reading low, moisture was in the dome. Signal cables were troubleshot on Feb 19, no problem found. It was assumed moisture was also in the body of instrument. Instrument was removed from service on Feb 20 and dried, however instrument still reads low. It was later discovered that Eppley AC ventilator fan had stalled and was creating a large amount of heat creating a large negative offset in signal. Until fan can be replaced the fuse was removed so that it does not generate heat.
3BMSDiffuse PSP12/28/201311:55:0012/28/201315:59:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, CUVA2, CUVB2, TUVR; and the following instruments were un-shaded: Diffuse 8-48, PSP, PIR. The South Sci-Tec tracker for these instruments are protected by UPS, however the battery depleted at 11:55 MST (the power outage started at 6:46 MST)
4BMSDiffuse PSP12/21/201310:30:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, CUVA2, CUVB2, TUVR; and the following instruments were un-shaded: Diffuse 8-48, PSP, PIR. The South Sci-Tec tracker for these instruments are protected by UPS, however the battery depleted at 10:30 MST (the power outage started at 7:00 MST)
5BMSDiffuse PSP09/06/201314:13:0009/06/201315:00:00SRRL BMS ventilators were cleaned on Sept 6, 2013. Due to prolonged un-shading of the Diffuse 8-48 and PSP, the data from 14:13 through 15:00 MST was corrected using data from the Diffuse CM22.
6BMSDiffuse PSP03/02/201313:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4. The Sci-Tec tracker for these insruments are protected by UPS, however the battery depleted at 13:06 MST (the power outage started at 8:06 MST)
7BMSDiffuse PSP05/07/201105:55:0005/09/201107:58:00There was a planned power outage for SRRL construction activities that started at 5:18 MST. Unfortunately, the backup generator provided to SRRL was providing bad power to the UPS powering the Sci-Tec trackers, therefore the UPS failed at 5:55 MST. Unfortunately, when normal AC power was restored at 11:45 MST, the Sci-Tec trackers had lost all their configuration due to UPS cycling on and off during the outage, therefore were not tracking. The tracker settings were restored on May 9 at 7:58 MST. The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4
8BMSDiffuse PSP05/27/201018:00:0005/28/201012:15:00Due to improper mounting of the replacement CH1 s/n 080033 the Sci-tec tracker was not tracking properly due to collision with the CH1 against tracker. The following BMS instruments were not pointing at the sun: Direct CH1, NIP2, CUVA2, & CUVB2 and the following instruments were unshaded: 8-48, PSP, PIR.
9BMSDiffuse PSP12/09/200912:11:0012/09/200914:22:00Due to a power outage at NREL, followed by bad power, the SRRL tracker UPS shutdown. The following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, NIP-RG780, LI201, TWC500, CUV4; and the following instruments were unshaded: 8-48, CM22, PSP, CUV4, PIR, CG4.
10BMSDiffuse PSP11/15/200814:02:0011/15/200814:32:00Due to a planned power outage at SRRL, the following BMS instruments were unshaded: Diffuse 8-48, CM22, PSP, CUV4, DW-PIR, DW-CG4.
11BMS17878F3Diffuse PSP07/07/200700:00:0007/09/200707:00:00The SRRL BMS Sci-Tec Tracker #1 shading arm was bent out of position. The PSP was partially unshaded during this time.
12BMS17878F3Diffuse PSP05/25/200610:50:0005/25/200611:10:00Sci-Tec tracker #1 tracker was down due to power shutdown for installation of new SRRL equipment lift. The following instruments were UNSHADED: Diffuse 8-48, Diffuse PSP, and the Downwelling PIR.
13BMS17878F3Diffuse PSP02/27/200609:16:0002/28/200607:57:00The SRRL BMS Diffuse Ventilated PSP was also removed from service, so that the instrument's RSnet could be characterized in the NREL Black Body System. It's output is set to missing (-99999).
14BMS17878F3Diffuse PSP07/23/200510:16:0007/23/200517:22:00Sci-Tec tracker #1 was down due to a scheduled NREL power outage on 7/23/2005 at 5:33 MST. The following instruments were UNSHADED: Diffuse 8-48, Diffuse PSP, and the Downwelling PIR.
15BMS17878F3Diffuse PSP07/10/200509:30:0007/10/200523:59:00The SRRL BMS Sci-Tec Tracker #1 was hung up on cables during this time. The following instruments were UNSHADED: Diffuse 8-48, Diffuse PSP and the PIR.
16BMSDiffuse PSP05/02/200313:45:0007/03/200308:30:00On May 2, 2003, the shade on the CM22 and PSP was removed during BORCAL, which caused the diffuse instruments to measure global. However on May 5, 2003, it was decided not to corrupt the baseline diffuse measurements with "global data", so the instruments were disconnected and/or shorted at the baseline datalogger connection. The instruments were later returned to normal service on July 3, 2003. All data during this time is bad or bogus.
17BMSDiffuse PSP02/26/200209:30:0004/01/200211:39:00Due to a shortage of diffuse trackers, the BMS Diffuse PSP was removed from service on Feb 26 @ 9:30 MST, to allow for another 8-48 to be calibrated in it's place. The instrument output was shorted on Feb 27 @ 7:47 MST to indicate that it is out of service. Note: This also affects the calculated Diffuse PSP (thermal corrected).
18BMSDiffuse PSP08/01/200000:00:0010/31/200023:59:00SRRL BMS was undergoing upgrade and reinstallation on the new instrument deck during this time. To keep a continuos data set for BMS, Diffuse PSP data from SRRL BSRN was used. Note: The following 25 days the BSRN diffuse tracker was malfunctioning: Aug 1, 2, 4, 7, 21; Sept 1-7, 12, 24-27; Oct 2, 8, 9, 12, 13, 15-17. The data for BMS diffuse was calculated on those days using global and direct normal. However both the Diffuse and Direct Trackers failed on Aug 5, 6, 17, 18. BMS Diffuse and Direct data on those 4 days were set to missing values (-99999).